HelpMePlease:

Simplify each expression. Use positive exponents. m^3 n^(–6) p^0

Champion:

any number to the power of zero is one

Champion:

and \[b^{-z}=\frac{1}{b^z}\] so \[n^{-6}=\frac{1}{n^6}\]